# Planet formation regulated by galactic-scale interstellar turbulence

### v2

@@ Abstract
Planet formation occurs over a few Myr within protoplanetary discs of dust and gas, which are often assumed to evolve in isolation. However, extended gaseous structures have been uncovered around many protoplanetary discs, suggestive of late-stage in-fall from the interstellar medium (ISM). To quantify the prevalence of late-stage in-fall, we apply an excursion set formalism to track the local density and relative velocity of the ISM over the disc lifetime. We then combine the theoretical Bondi-Hoyle-Lyttleton (BHL) accretion rate with a simple disc evolution model, anchoring stellar accretion time-scales to observational constraints. Disc lifetimes, as well as masses, stellar accretion rates and approximate gas radii as a function of stellar mass and age are remarkably well-reproduced by our simple model that includes only ISM accretion. If BHL accretion is the origin of the initial disc material, then $20{-}70$ percent of discs are mostly composed of material accreted in the most recent half of their lifetime, suggesting that disc properties are not a direct test of isolated evolution models. The ISM also deposits sufficient energy to drive observed turbulence in protoplanetary discs with viscous $\alpha_{\mathrm{SS}}\sim 10^{-5}-10^{-1}$, although we emphasise that angular momentum transport may still be driven internal processes. Overall, our findings represent a compelling case for regulation of planet formation by galactic-scale turbulence, with broad consequences for planet formation theory. Our simple approach can be easily applied to semi-analytic models, and our results represent urgent motivation for an unbiased survey to confirm or refute our findings.
@@ 1 Introduction
Recent observations in molecular line emission (Huang et al., 2021) and infrared scattered light (Garufi et al., 2020; Benisty et al., 2023) are frequently revealing extended structures associated with protoplanetary discs, on spatial scales several times the disc outer radius. In-fall has been suggested as a possible origin for these structures (Dullemond et al., 2019; Kuffmeier et al., 2020, 2023; Hanawa et al., 2024), which are highly complex, including large scale spirals and streamers around T Tauri (Huang et al., 2023) and Herbig AeBe (Boccaletti et al., 2020) stars, often associated with reflection nebulosity (Gupta et al., 2023), accretion outbursts (Pérez et al., 2020; Hales et al., 2020; Dong et al., 2022) and misaligned inner discs (Ginski et al., 2021). Evidence that is suggestive of in-fall events has been found not only during the earliest stages of disc evolution, when the star is still embedded in its natal environment (Jørgensen et al., 2009; Maury et al., 2019; Codella et al., 2024), but also around stars such as AB Aur (Nakajima & Golimowski, 1995; Grady et al., 1999; Fukagawa et al., 2004) DR Tau (Mesa et al., 2022) and SU Aur (Ginski et al., 2021), which are all mature star-disc systems of age $\sim 1{-}3$ Myr. A recent estimate from a sample size of $43$ disc-hosting stars suggests approximately $16$ percent are interacting with the ambient ISM (Garufi et al., 2024). This material should fall onto the disc on a time-scale much shorter than the disc lifetime (Dullemond et al., 2019; Kuffmeier et al., 2020), suggesting that either these events occur much more frequently than observed or the external structure is replenished by the ISM for a prolonged period.
@@
In-falling gas may be the consequence of Bondi-Hoyle-Lyttleton (BHL) accretion, where low velocity gas is captured due to the gravitational potential of the star (Bondi, 1952; Edgar, 2004). Theoretically, the rate of capture of material from the interstellar medium (ISM) via BHL accretion can be substantial while stars remain bound to a dense gas reservoir. Padoan et al. (2005) proposed this mechanism to explain the observed relationship between stellar accretion rates and stellar mass. This idea was supported by subsequent work (Throop & Bally, 2008; Klessen & Hennebelle, 2010), while more recent analysis of turbulent hydrodynamic simulations in a periodic box have shown that late stage accretion may be significant source of mass and angular momentum for $\sim 1$ Myr old discs (Kuffmeier et al., 2023; Pelkonen et al., 2024; Padoan et al., 2024). However, assessing the importance of this process for the observed population of protoplanetary discs throughout their lifetime requires quantifying the temporal evolution of the turbulent ISM surrounding forming stars. This is challenging for hydrodynamic simulations that are both expensive and temporally limited by the turbulent crossing time of the computational domain.
@@
In this work, we present a semi-analytic approach for tracking the BHL accretion rate over the disc lifetime. In a sense, this is a complement and extension of that of Throop & Bally (2008), who tracked BHL accretion rates onto stars that are gravitationally bound to a dense gas reservoir. Our aim here is to estimate the distribution of accretion histories over the whole disc lifetime for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. This includes both the period in which the star is bound to the gaseous overdensity in which it forms (Throop & Bally, 2008) and once the bound gas has dispersed. Statistically interpreting the role of BHL accretion on a disc population level requires simulating the turbulent cascade from galactic to protoplanetary disc size scales for the several Myr over which discs persist (Klessen & Hennebelle, 2010; Colman et al., 2022; Nusser & Silk, 2022). To do so, we apply an excursion set formalism (Hopkins, 2012) which is effectively zero dimensional. This is therefore a highly efficient way of statistically following density fluctuations of the ISM, making it conducive to semi-analytic planet formation models (e.g. Weder et al., 2023; Mordasini & Burn, 2024).
We adopt the following approach:
  1. 1.
    Properties of local star forming regions: Apply the excursion set formalism to determine the typical properties of star forming regions that are targeted for local disc surveys (within $\sim 200$ pc).
  2. 2.
    Collapse of the molecular cloud: Adopt a simple model for the collapse of these star forming regions, and the rate at which stars form as a function of time.
  3. 3.
    Stellar dynamics in the bound cloud: Draw a representative stellar population, using the star formation rate based on step 2 to statistically calculate when stars should form during the collapse of a region. Compute the dynamical evolution of these stars under the cloud potential until cloud dispersal, tracking the local gas density and velocity.
  4. 4.
    Evolution of the turbulent medium: Compute the time-evolution of the turbulent medium outside of the bound star forming region, using the same approach as in step 1, to continue to track the local density and velocity after the exhaustion of the bound gas reservoir (star forming region).
  5. 5.
    Accretion rate onto the star-disc system: Calculate the BHL accretion rate onto the star-disc system over time.
  6. 6.
    Disc properties: Couple the external accretion rate to a simple model including accretion onto the host star, to track the evolution of the disc masses, radii, turbulence and stellar accretion rate.
  7. 7.
    Observational comparison: Compare the resultant model to observed disc properties.
In the remainder of this work, we describe these steps in greater detail, and discuss the implications for protoplanetary disc evolution.
@@ 2.1 Overview
In this section, our aim is to estimate the distribution of accretion histories for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. To do so, we must follow the velocity and density distribution of gas surrounding these young stars. This includes both the period in which it is bound to the gaseous overdensity in which it forms (cf. Throop & Bally, 2008), and once the bound gas has dispersed. We estimate the statistical evolution of the accretion rate semi-analytically by applying an excursion set formalism (Hopkins, 2012). This method is effectively ‘0D’, and is therefore a highly efficient way of statistically following density fluctuations of the ISM.
We adopt the following approach in the remainder of this section. First, we introduce the excursion set formalism and apply it to review the typical properties of local giant molecular clouds (GMCs) and star forming regions (Section 2.2). We then introduce a simple model for star formation during the collapse of bound GMCs (Section 2.3), and the approach for modelling the dynamics of stars forming in this environment (Section 2.4). We consider the temporal evolution of density and velocity of the unbound, galactic ISM outside of the star forming region in Section 2.5. Finally, we evaluate how the accretion rate for individual stars evolves as they form from the local population of GMCs (Section 2.6).
@@ 2.2 Properties of local star forming regions
Figure 1: Sampling of the mass function via the excursion set formalism for GMCs (blue points) based on $10^{5}$ randomly drawn density perturbation trajectories in the solar neighbourhood. We have normalised the $y$-axis such that we show the average number of regions younger than $10$ Myr that are expected to be found within $200$ pc. We highlight in yellow the region which is well-sampled, corresponding to typical local star forming regions which are surveyed for protoplanetary disc properties. In cyan, we show the range of initial GMC masses we consider to generate our disc population synthesis. In orange, we show the same mass function assuming a fixed star formation efficiency $\epsilon=0.5$.
In the first instance, we wish to quantify the properties of GMCs that undergo gravitational collapse to form stars. This can be reframed in terms of quantifying the maximum turbulent size-scale on which a given volume of space in the mid-plane of the galaxy is bound at any given time. All scales smaller than this size-scale will also be bound to to this cloud, and therefore this can be considered a unit of star formation, or a star forming region. This problem can be understood as a barrier-crossing problem arising from a random walk through Fourier space, which can be solved using excursion set theory as has been successfully applied to dark matter halos (Bond et al., 1991).
We follow Hopkins (2012) who applied the excursion set formalism in the context of GMCs in the galactic disc. The approach can be summarised as follows, but we refer the interested reader to that work for greater detail. Most generally, it is clear that the density of gas on any given scale can be described as the sum of the contributions to this density at all larger scales. Supersonic turbulence yields a log-normal probability distribution in volume space for the gas density, with a dispersion that is dependent on the local Mach number (Vazquez-Semadeni, 1994). Therefore, dispersion in the contribution of any range of scales to the deviation from the median depends on the turbulent velocity on that scale. In a supersonic, rapidly cooling medium, the cascade of energy from large scales to small scales produces an energy spectrum with power-law exponent $p=2$, or turbulent velocity that scales with the square-root of the size-scale (Burgers, 1948). The critical density required for collapse at any intermediate size-scale is then related to the effective pressure support afforded to the medium from the turbulent motions on that scale, via the dispersion relation for a disc of a finite thickness (Begelman & Shlosman, 2009). On the largest scales, this pressure support effectively comes from galactic shear, while on the smallest scales, turbulent fragmentation can be suppressed by the thermal pressure (Jeans, 1902). In principle, the influence of magnetic fields can also be accounted for in this context, but Hopkins (2012) notes that this modification practically amounts to changing the sound speed, with only minor quantitative consequences for our model. In the context of the BHL accretion rate that we eventually compute, modifications of only a factor of order unity are expected due to local magnetic fields (Lee et al., 2014; Burleigh et al., 2017). Given the turbulent energy spectrum, a ‘trajectory’ in logarithmic density space can be instantaneously defined as the cumulative sum of perturbations from the largest to smallest scales. For any trajectory, a gravitationally unstable GMC, or star forming region, is defined on the largest length scale which is unstable (if any), all smaller length scales being subsumed within the larger volume.
In our case, we follow similar parameter choices to those adopted by the fiducial case of Hopkins (2012). The energy spectrum normalisation is defined such that the velocity dispersion at the scale height $h$ of the gaseous galactic disc $\sigma_{\rm g}(h)=h\Omega=6$ km s-1, where $\Omega=2.6\times 10^{-2}$ Myr-1. We assume a surface density of the galactic disc in the solar neighbourhood $\Sigma=12\,M_{\odot}$ pc-2, yielding a Toomre $Q=1.33$. We always assume a sound speed $c_{\mathrm{s}}=0.2$ km s-1 and define a grid with a maximum scale $R_{\mathrm{max}}=5h$ down to a minimum scale $R_{\mathrm{min}}<0.01$ pc, each grid cell being $95$ percent of the spatial scale of the previous cell.
Finally, to define the absolute frequency of star forming regions within a certain age range in the solar neighbourhood depends on the turbulent time-scale. This is itself a function of spatial scale:
@@
where $t_{\rm cross}$ is the crossing time-scale and $\eta\approx 1$ (Pan & Scannapieco, 2010). An appropriate time-scale ${\tau}_{\mathrm{t,0}}$ over which the trajectory ‘resets’ corresponds to the turbulent time-scale on the scales that dominate density variations. We therefore estimate ${\tau}_{\mathrm{t,0}}\approx\tau_{\mathrm{t}}(h)$. We can then estimate the differential number of star forming region younger than a certain age $\Delta\tau$ within a given volume $\Delta V$:
$\left\langle\frac{\mathrm{d}N}{\mathrm{d}M}\right\rangle\approx\frac{1}{2^{3}}\frac{\Delta\tau}{\tau_{\mathrm{t,0}}}\frac{\Delta V\rho}{M}\cdot\frac{N_{\mathrm{coll}}}{N_{\mathrm{traj}}}$ (2)
where $N_{\mathrm{coll}}$ is the number of unstable trajectories from a sample of $N_{\mathrm{traj}}$. We include a factor $1/2^{3}$ since given a region of radius $R_{\mathrm{SFR}}$, any other region with $R_{\mathrm{SFR}}$ is forbidden inside a distance of $2R_{\mathrm{SFR}}$ (such a region would merge with the original region). In the averaging for equation 2 we have ignored that similar mass unstable regions may be spatially correlated, since the probability of collapse on a given spatial scale is dependent on the density contribution from larger scales. However, Hopkins (2012) find that this bias is negligible for the low mass star forming regions, with $R_{\mathrm{SFR}}\ll h$. The shape of the mass function for the local, low mass GMCs relevant in the context of this work should therefore be correct. Clearly, star forming regions should still preferentially be found where the gas density is highest, we assume that using the average surface density for a sufficient $R_{\mathrm{obs}}\gtrsim h$.

### v3

@@ Abstract
Planet formation occurs over a few Myr within protoplanetary discs of dust and gas, which are often assumed to evolve in isolation. However, extended gaseous structures have been uncovered around many protoplanetary discs, suggestive of late-stage in-fall from the interstellar medium (ISM). To quantify the prevalence of late-stage in-fall, we apply an excursion set formalism to track the local density and relative velocity of the ISM over the disc lifetime. We then combine the theoretical Bondi-Hoyle-Lyttleton (BHL) accretion rate with a simple disc evolution model, anchoring stellar accretion time-scales to observational constraints. Disc lifetimes, masses, stellar accretion rates and gaseous outer radii as a function of stellar mass and age are remarkably well-reproduced by our simple model that includes only ISM accretion. We estimate $20{-}70$ percent of discs may be mostly composed of material accreted in the most recent half of their lifetime, suggesting disc properties are not a direct test of isolated evolution models. Our calculations indicate that BHL accretion can also supply sufficient energy to drive turbulence in the outer regions of protoplanetary discs with viscous $\alpha_{\mathrm{SS}}\sim 10^{-5}-10^{-1}$, although we emphasise that angular momentum transport and particularly accretion onto the star may still be driven by internal processes. Our simple approach can be easily applied to semi-analytic models. Our results represent a compelling case for regulation of planet formation by large-scale turbulence, with broad consequences for planet formation theory. This possibility urgently motivates deep observational surveys to confirm or refute our findings.
@@ 1 Introduction
Recent observations in molecular line emission (Huang et al., 2021) and infrared scattered light (Garufi et al., 2020; Benisty et al., 2023) are frequently revealing extended structures associated with protoplanetary discs, on spatial scales several times the disc outer radius. In-fall has been suggested as a possible origin for these structures (Dullemond et al., 2019; Kuffmeier et al., 2020, 2023; Hanawa et al., 2024), which are highly complex, including large scale spirals and streamers around T Tauri (Huang et al., 2023) and Herbig AeBe (Boccaletti et al., 2020) stars, often associated with reflection nebulosity (Gupta et al., 2023), accretion outbursts (Pérez et al., 2020; Hales et al., 2020; Dong et al., 2022) and misaligned inner discs (Ginski et al., 2021). Evidence that is suggestive of in-fall events has been found not only during the earliest stages of disc evolution, when the star is still embedded in its natal environment (Jørgensen et al., 2009; Maury et al., 2019; Codella et al., 2024), but also around stars such as AB Aur (Nakajima & Golimowski, 1995; Grady et al., 1999; Fukagawa et al., 2004) DR Tau (Mesa et al., 2022) and SU Aur (Ginski et al., 2021), which are all mature star-disc systems of age $\sim 1{-}3$ Myr. A recent estimate from a sample size of $43$ disc-hosting stars suggests approximately $16$ percent are interacting with the ambient ISM (Garufi et al., 2024). This material should fall onto the disc on a time-scale much shorter than the disc lifetime (Dullemond et al., 2019; Kuffmeier et al., 2020), suggesting that either these events occur much more frequently than observed or the external structure is replenished by the ISM for a prolonged period. Indeed, substantial mass accumulation at least over the first $\sim 1$ Myr of the disc life-time is hinted at by the fact that discs younger than $1$ Myr in $\rho$ Ophiucus and Corona Australis are somewhat more compact and lower mass than observed in $\sim 1{-}2$ Myr old regions (Testi et al., 2016, 2022; Cieza et al., 2019; Cazzoletti et al., 2019; Williams et al., 2019). It is possible that discs contain a substantial fraction of recently accreted material for much of their lifetime, which would have significant consequences for disc evolution models.
@@
In-falling gas may be the consequence of Bondi-Hoyle-Lyttleton (BHL) accretion, where low velocity gas is captured due to the gravitational potential of the star (Bondi, 1952; Edgar, 2004). Theoretically, the rate of capture of material from the interstellar medium (ISM) via BHL accretion can be substantial while stars remain bound to a dense gas reservoir. Padoan et al. (2005) proposed this mechanism to explain the observed relationship between stellar accretion rates and stellar mass. This idea was supported by subsequent work (Throop & Bally, 2008; Klessen & Hennebelle, 2010), while more recent analysis of turbulent hydrodynamic simulations in a periodic box have shown that late stage accretion may be significant source of mass and angular momentum for $\sim 1$ Myr old discs in high gas density environments (Kuffmeier et al., 2023; Pelkonen et al., 2024; Padoan et al., 2024). In-fall has also been suggested as a mechanism to explain examples of substantial protoplanetary discs persisting around old stars (Scicluna et al., 2014; Derkink et al., 2024). However, assessing the importance of this process for the observed population of protoplanetary discs throughout their lifetime requires quantifying the temporal evolution of the turbulent ISM surrounding forming stars. This is challenging for hydrodynamic simulations that are both expensive and limited in integration time by the (turbulent) crossing time of the computational domain.
@@
In this work, we present a semi-analytic approach for tracking the BHL accretion rate over the disc lifetime. In a sense, this is a complement and extension of that of Throop & Bally (2008), who tracked BHL accretion rates onto stars that are gravitationally bound to a dense gas reservoir. Our aim here is to estimate the distribution of accretion histories over the whole disc lifetime for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. This includes both the period in which the star is bound to the gaseous overdensity in which it forms (Throop & Bally, 2008) and once the bound gas has dispersed. Statistically interpreting the role of BHL accretion on a disc population level requires simulating the turbulent cascade from galactic to protoplanetary disc size scales for the several Myr over which discs persist (Klessen & Hennebelle, 2010; Colman et al., 2022; Nusser & Silk, 2022). To do so, we apply an excursion set formalism which is effectively zero dimensional (Hopkins, 2012). This is therefore a highly efficient way of statistically following density fluctuations of the ISM, making it conducive to semi-analytic planet formation models (e.g. Weder et al., 2023; Mordasini & Burn, 2024).
In the remainder of this work, we first present a model for tracking the local ISM conditions and resultant BHL accretion rate for stars evolving in a turbulent ISM (Section 2). We then present a simple disc model and compare to observed disc properties in Section 3. We discuss our results in the broader context of planet formation in Section 4 and draw conclusions in Section 5.
@@ 2.1 Overview
In this section, our aim is to estimate the distribution of accretion histories for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. To do so, we must follow the velocity and density distribution of gas surrounding these young stars. This includes both the period in which the star remains within the star forming region (cf. Throop & Bally, 2008), and once the bound giant molecular gas has dispersed. The latter, and to some degree the former, requires quantifying the properties of the supersonic, turbulent ISM.
Simulations of sustained, supersonic turbulent flows that are not self-gravitating have clearly shown that the volume-weighted probability density function for density fluctuations is lognormal, with a variance that is a weak function of the Mach number (e.g. Vazquez-Semadeni, 1994; Ostriker et al., 1999). A small fraction of the ISM in a steady state is therefore at a sufficient density to marginally overcome (turbulent) pressure support, and undergo gravitational collapse (Evans, 1999). Regions which meet these criterion are known as giant molecular clouds (GMCs). These structures have long been known to obey so-called Larson relations, which broadly connect the size scale $R_{\mathrm{cloud}}$ and velocity dispersion $\sigma_{v}$ of molecular clouds in the galactic ISM (Larson, 1981). Since original relations, improved data has established that generally the velocity dispersion $\sigma_{v}\propto R_{\mathrm{cloud}}^{1/2}$ (Heyer & Brunt, 2004), as expected from supersonic, rapidly cooling turbulence (Burgers, 1974), although the normalisation of this relation may somewhat vary with environment (Green et al., 2024). Any models aiming to track turbulent fluctuations in the ISM density and velocity, or the resultant properties of collapsing GMCs, must take into account these relationships.
To track turbulent fluctuations in the ISM, we adopt the excursion set approach described by Hopkins (2012), with the key equations summarised in Appendix A for convenience. In brief, this approach involves quantifying the statistical distribution of densities or velocities at a given location in space as the sum of contributions to that density on larger scales (Appendix A.1). A ‘trajectory’ is defined as the cumulative sum of perturbations from the largest scale to the smallest scales (equation A9). These trajectories define the density and velocity on a single volume element, although the size of the volume element depends on the scale being considered. For any given trajectory, we can determine whether the local density is sufficient to undergo gravitational collapse (Appendix A.2), becoming a star forming region. Each scale also has an associated turbulent time-scale, based on which we can compute the temporal evolution of the turbulent fluctuations (Appendix A.3). These latter two aspects allow us to estimate local density and velocity fluctuations during and after the star formation process.
In the remainder of this section, we explain the following steps in our procedure:
  1. 1.
    Quantify the properties of local star forming regions (Section 2.2): Apply the excursion set formalism to establish a statistically representative sample of unstable trajectories, and therefore determine the typical properties of star forming regions that are targeted for local disc surveys (within $\sim 200$ pc).
  2. 2.
    Follow the collapse of molecular clouds and star formation (Section 2.3): For a representative sample of star forming regions, with associated trajectories, we adopt a simple model for collapse and the star formation rate. This allows us to draw the formation times for a representative population of stars. Note that a star forming region may contain many stars, and is therefore not equivalent to the protostellar collapse or class 0 phase of stellar evolution.
  3. 3.
    Track the dynamical evolution of stars in the bound cloud (Section 2.4): For the representative stellar population, we track the dynamical evolution of stars based on the gravitational potential of the host star forming region.
  4. 4.
    Evolve temporal fluctuations of the turbulent medium (Section 2.5): Once the cloud is exhausted, the star is ejected into the turbulent medium. We continue to evolve the trajectory to track both density and velocity of the ISM during and after the collapse of the cloud according to equation A13.
  5. 5.
    Compute accretion rate onto the star-disc system (Section 2.6): Based on the local density and velocity during and after the collapse of the star forming region, we calculate the BHL accretion rate onto the star-disc system over time.
In the remainder of this section, we describe these steps in more detail. For each stage, we also highlight caveats to our approach, which amount to points for development in future work.
@@ 2.2 Properties of local star forming regions
Figure 1: Sampling of the mass function via the excursion set formalism for GMCs (blue points) based on $10^{5}$ randomly drawn density perturbation trajectories in the solar neighbourhood. Sampling errors are smaller than the marker size. We have normalised the $y$-axis such that we show the average number of regions younger than $10$ Myr that are expected to be found within $200$ pc. We highlight in yellow the region which is well-sampled, corresponding to typical local star forming regions which are surveyed for protoplanetary disc properties. In cyan, we show the range of initial GMC masses we consider to generate our disc population synthesis. In orange, we show the same mass function assuming a fixed star formation efficiency $\epsilon=0.5$.
2.2.1 Approach summary
We first need to quantify the mass function of gravitationally unstable GMCs. To do so, we follow the approach of Hopkins (2012), as described in Section 3 of that work. In brief, this involves computing the probability of the turbulent medium collapsing at a given scale, based on the turbulent energy spectrum. We randomly sample a number of ‘trajectories’, $N_{\mathrm{traj}}$, of which some number $N_{\mathrm{coll}}$ are unstable at some length scale $R$ corresponding to total mass $M$ (see Appendix A.2). When weighted by the inverse of the relative volume that this star forming region represents, this yields an estimate for the mass function of unstable GMCs. However, to estimate the statistically sampled upper limit of the mass function among local star forming regions we must also estimate a normalisation constant. The absolute frequency of star forming regions within a certain age range in the solar neighbourhood depends on the turbulent time-scale. This is itself a function of spatial scale:
@@ 2.2.1 Approach summary
where $t_{\rm cross}$ is the crossing time-scale and $\eta\approx 1$ (Pan & Scannapieco, 2010). The root mean square turbulent velocity $\langle v_{\rm t}^{2}(R)\rangle^{1/2}$ depends on the turbulent energy spectrum, as described in Appendix A. An appropriate time-scale ${\tau}_{\mathrm{t,0}}$ over which the trajectory ‘resets’ corresponds to the turbulent time-scale at given scale. For simplicity, here we adopt ${\tau}_{\mathrm{t,0}}\approx\tau_{\mathrm{t}}(h)$, reasoning that fluctuations at a scale $R\sim h$ are approximately the highest amplitude (we discuss this simplification in Section 2.2.3). We can then estimate the differential number $\mathrm{d}N$ of star forming region younger than a certain age $\Delta\tau$ within a given volume $\Delta V$:
$\left\langle\frac{\mathrm{d}N}{\mathrm{d}M}\right\rangle\approx\frac{1}{2^{3}}\frac{\Delta\tau}{\tau_{\mathrm{t,0}}}\frac{\Delta V\rho}{M}\cdot\frac{N_{\mathrm{coll}}}{N_{\mathrm{traj}}},$ (2)
where $\rho$ and $M$ are the density and mass of the star forming region respectively. We include a factor $1/2^{3}$ since given a region of radius $R_{\mathrm{SFR}}$, any other region with $R_{\mathrm{SFR}}$ is forbidden inside a distance of $2R_{\mathrm{SFR}}$ (such a region would merge with the original region).
2.2.2 Mass function